# Generating Model Parameters for Controlling: Parameter Diffusion for Controllable Multi-Task Recommendation

### v1

@@ 3.3 Parameter Diffusion Model Training
$\ell_{\text{diff}}=\mathbb{E}_{\bm{\theta}_{i,0}^{\text{a}},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\bm{\xi}}(\bm{\theta}_{i,t}^{\text{a}},\bm{w}_{i},t)\right\|^{2}\right],$ (8)
where $\epsilon$ denotes the noise to obtain $\bm{\theta}_{i,t}^{\text{a}}$ from $\bm{\theta}_{i,0}^{\text{a}}$, and the denoising model $\epsilon(\cdot)$ is the main part of the generative model $g_{\bm{\xi}}$. We assume that the parameters of $g_{\bm{\xi}}$ primarily originate from the denoising model. For simplicity, we denote the denoising model as $\epsilon_{\bm{\xi}}$. To conduct condition training in a classifier-free guidance manner (Ho & Salimans, 2022), we use the denoising model to serve as both the conditional and unconditional model by simply inputting a null token $\varnothing$ as the condition (i.e., preference weights $\bm{w}_{i}$) for the unconditional model, i.e. $\epsilon_{\bm{\xi}}(\bm{\theta}_{i,t}^{\text{a}},t)=\epsilon_{\bm{\xi}}(\bm{\theta}_{i,t}^{\text{a}},\bm{w}_{i}=\varnothing,t)$. The probability of setting $\bm{w}_{i}$ to $\varnothing$ is denoted as $p_{\mathrm{uncond}}$ and is configured as a hyperparameter. Alg. LABEL:alg:train in Appendix illustrates the detailed procedure.
3.4 Test-Time Parameter Generation
@@
$\tilde{\epsilon}_{\bm{\xi}}(\bm{\theta}_{n,t}^{\text{a}},\bm{w}_{n},t)$ $=(1+\gamma)\epsilon_{\bm{\xi}}(\bm{\theta}_{n,t}^{\text{a}},\bm{w}_{n},t)-\gamma\epsilon_{\bm{\xi}}(\bm{\theta}_{n,t}^{\text{a}},t),$ (9)
@@
where $\bm{z}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ for $t>1$ and $\bm{z}_{t}=\bm{0}$ for $t=1$, $\beta_{t}=1-\alpha_{t}$, $\gamma\in[0,1]$ Alg. LABEL:alg:sample in Appendix illustrates the detailed procedure.
Figure 2: The relationship between the loss values in Eq. (7) and the adapter parameters $\bm{\theta}_{i}^{\text{a}}$ (with the preference weights $w_{i}^{1}=0.3$ and $w_{i}^{2}=0.7$) on MovieLens 1M using SASRec as backbone. The index order indicates the number of epochs after convergence. The detailed settings are shown in Sec. LABEL:sec:appendix:discussion
@@
4 Discussions on the robustness of parameter diffusion
In experiments of Sec. 5, we found that parameter diffusion in our PaDiRec can generate model parameters that differ from those obtained by retraining the total loss in Eq. (7), yet still achieve good recommendation performance and controllability. This leads us to hypothesize that parameter diffusion may find more robust model parameters through model parameter generalization. To validate this hypothesis, we continued training the adapter parameters for multiple epochs after the total loss had converged, analyzing the relationship between the adapter parameters and the values of the loss Eq. (7). We employed polynomial fitting to model the relationship between the data points, estimating the functional relationship between the loss values and model parameters, achieving a goodness of fit of $R^{2}=0.74$. As illustrated in Figure 2, we observe that the solutions of the total loss exhibit different characteristics under various model parameters. Specifically, there are relatively flat solution sets, referred to as “Stable Solutions”, as well as solution sets with more dramatic fluctuations, termed “Sharp Solutions”. Furthermore, as verified in RQ1 of Sec. 5.3, we observe that the parameter diffusion in PaDiRec effectively learns the set of robust “Stable Solutions”, thereby enhancing controllability while maintaining high performance.
5 Experiments
We conducted experiments to evaluate the performance of PaDiRec on two public datasets and an industrial dataset for sequential recommendation.
5.1 Experiment settings
Dataset. We used two public datasets, MovieLens 1M  [^1] https://grouplens.org/datasets/movielens/ and Amazon Food  [^2] http://jmcauley.ucsd.edu/data/amazon/links.html, and the Industrial Data from a electronics commercial store. Detailed descriptions of the datasets and preprocessing methods can be found in Appendix LABEL:sec:dataset.
Baselines. The baselines are as follows. Retraining as Eq. (2), which is considered optimal. Soup (Wortsman et al., 2022), a classic algorithm for model merging. MMR (Carbonell & Goldstein, 1998), a rule-based post-process policy. CMR (Chen et al., 2023), an re-rank algorighm utilizing hypernetwork to achieve dynamic preference of changing. LLM (Appendix  LABEL:sec:llm:prompt), a prompt-based method for controllable recommendation. Details about all the baselines are shown in Appendix LABEL:sec:appendix:baselines
@@ 5.1 Experiment settings
Metrics. We propose evaluating performance from two dimensions. Specifically, we use Hypervolume (HV) (Guerreiro et al., 2021) to measure the performance of the algorithm on each task, particularly in terms of the trade-offs between accuracy and diversity. The average HV (denoted as Avg.HV) across multiple tasks (i.e., various preference weights) is used to assess the overall performance of the algorithm in balancing both objectives (accuracy and diversity). To eliminate the differences in scale between the two objectives, we normalize the performance on each objective. Additionally, we utilize the Pearson correlation coefficient to evaluate the alignment between the algorithm’s performance across different tasks and the optimal model, providing insight into the algorithm’s controllability. Pearson r-a, Pearson r-d measure the correlation between the algorithm and the optimal in terms of accuracy (NDCG@$10$) and diversity ($\alpha$-NDCG@10).
5.2 Experimental results
We conducted experiments to address the following two questions: i) How transferable is PaDiRec, specifically in terms of its ability to adapt to different backbone algorithms? ii) How does PaDiRec perform compared to other baselines on each specific task? The results are presented in Table LABEL:tab:main.
To answer the first question, we used commonly adopted sequential recommendation models as backbones (e.g., SASRec (Kang & McAuley, 2018), GRU4Rec (Hidasi, 2015), and TiSASRec (Li et al., 2020b)) and conducted extensive experiments across three datasets. Specifically, we evaluated PaDiRec’s performance under various task descriptions by measuring NDCG@10 and $\alpha$-NDCG@10. The accuracy weight $w_{\text{acc.}}$ varies from 0 to 1 in intervals of 0.1, with the corresponding diversity weight set as $w_{\text{div.}}=1-w_{\text{acc.}}$. We then post-processed NDCG@10 and $\alpha$-NDCG@10 across different tasks to compute Avg.HV, Pearson r-a, and Pearson r-d. These metrics respectively evaluate the quality of multi-objective optimization on individual tasks and the controllability across multiple tasks. Overall, across the three backbones and three datasets, PaDiRec consistently ranked among the top two performers across all three evaluation metrics. Notably, in most cases, the top two of Avg.HV are Retrain and PaDiRec, indicating that PaDiRec’s performance in multi-objective trade-offs is on par with, or even superior to, Retrain. Specific exceptions occurred, such as on the Amazon Food dataset with GRU4Rec as the backbone and the Industrial Data with TiSASRec as the backbone, where CMR achieved the best Avg.HV. This is because CMR is not influenced by task descriptions and thus maintains consistently high NDCG@10 scores (as shown in Figure 3 and further explained in response to question ii). For Pearson r-a and Pearson r-d, PaDiRec demonstrated strong correlations with the Retrain method, indicating that PaDiRec closely aligns with Retrain (which we assume to be optima) in terms of accuracy (NDCG@10) and diversity ($\alpha$-NDCG@10) across different tasks. On the Amazon Food dataset with SASRec as the backbone, CMR achieved the highest Pearson r-d. However, its Pearson r-a was negative, indicating a lack of control and a collapse in accuracy.
@@ 5.2 Experimental results
(b) Accuracy on Amazon Food
(c) Accuracy on Industrial Data
(d) Diversity on MovieLens 1M
(e) Diversity on Amazon Food
@@
Figure 3: The accuracy and diversity curve of PaDiRec and other baselines in NDCG@10 and $\alpha$-NDCG@10 across accuracy weights ranging from 0 to 1, with intervals of 0.1. The backbone is TiSASRec, results on other backbones are shown in Appendix LABEL:sec:appendix:tend
To address the second question, we presented the specific performance of PaDiRec under each task description using TiSASRec as the backbone across three datasets, as shown in Figure 3 (more results are shown in Appendix. LABEL:sec:appendix:tend). It is observed that in all three datasets, PaDiRec’s NDCG@10 progressively increases with higher accuracy weights, while $\alpha$-NDCG@10 decreases correspondingly due to the simultaneous reduction in diversity weight. These trends demonstrate the effectiveness of our algorithm in controllability. Notably, assuming that Retrain is optimal, PaDiRec exhibits strong consistency with the Retrain method. In contrast, MMR, as a post-processing algorithm, shows variability because varying degrees of diversity manipulation can disrupt the original recommendation list, uncontrollably affecting its accuracy. The Soup method merges the parameters of accuracy and diversity models based on their weights, aligning closely with the Retrain model when accuracy weights are extreme but showing significant deviations in other tasks. This indicates that tasks do not follow a simple linear relationship with different preference weights, and Soup makes overly strong assumptions about this relationship. CMR demonstrates inconsistent performance across different datasets. On the MovieLens 1M dataset, CMR aligns well with the original descriptions by exhibiting high diversity. However, on the other two datasets, it shows a stable yet uncontrollable state; for instance, on the Amazon Food dataset, CMR maintains high accuracy even with low accuracy weights, and on the Industrial Data, it retains high diversity despite low diversity weights.
5.3 Analyses
We conducted our analysis experiments based on three key research questions: RQ1: What are the advantages of Diffusion over Hypernetwork in parameter generation? RQ2: Is PaDiRec efficient enough to handle real-time changes in preference weights compared to Retrain? RQ3: How do different conditioning strategies impact the model’s performance? Additionally, we also present some case study in Apprndix LABEL:sec:appendix:case.
Regarding RQ1: Diffusion outperforms Hypernetwork in parameter generation.
We conducted experiments to validate the robustness of the parameters generated by PaDiRec (as assumed in Sec. 4). Specifically, we designed three sets of experiments to constructed adapter parameters: “Retrain”, “PaDiRec” and “Hypernetwork”, where “Hypernetwork” utilizes the MLP to learn the relationship between the preference weight and the optimized adapter parameters.
@@ 5.3 Analyses
Figure 4: The variation of performance before and after disturbance in MovieLens 1M based on SASRec. The blue bars represent the variation in NDCG@10, while the red represent the variation in $\alpha$-NDCG@10
First, we added Gaussian noise of the same magnitude to all three sets of adapter parameters and measured the resulting fluctuations in NDCG@10 and $\alpha$-NDCG@10. The experiments were repeated multiple times under various preference weights, and the results are shown in Figure 4. We observed that the parameters generated by PaDiRec exhibited the lowest performance fluctuations, both in terms of accuracy (NDCG@10) and diversity ($\alpha$-NDCG@10), when subjected to perturbations, which verifies the assumption in Sec. 4. Additionally, we compared the similarity (Inverse Euclidean Distance) between the three sets of parameters and the Retrain parameters. Hypernetwork-generated parameters showed slightly higher similarity to the Retrain parameters than those generated by Diffusion, with values of $0.3147\pm 0.0184$ and $0.3035\pm 0.0184$, respectively. This indicates that PaDiRec, as a diffusion-based parameter generator, is not merely mimicking parameters but has learned the underlying distribution of parameters, demonstrating its ability to generate robust, high-performance parameters (Wang et al., 2024).
Regarding RQ2: The efficiency and effectiveness of PaDiRec.
PaDiRec is designed to adaptively adjust model parameters in an online environment without retraining, enabling it to quickly respond to new task requirements. This places a strong emphasis on the model’s response time. We compared the response times of PaDiRec and “Retrain” across various backbones and datasets, with the results shown in Table LABEL:tab:time. As observed, in all experiments using three different backbones across three datasets, PaDiRec’s response time was significantly faster than that of “Retrain”. Notably, the response time of “Retrain” correlated with the size of the dataset, whereas PaDiRec exhibited minimal variation across different datasets. This highlights PaDiRec’s data-agnostic nature indicating its potential for handling large-scale datasets efficiently.
Regarding RQ3: Influence of different conditioning strategies.
We investigate the influence of different conditioning strategies aimed at improving the integration of conditions into the denoising model. As shown in Figure 5, each strategy emphasizes different performance dimensions (details on the construction of each strategy can be found in Appendix LABEL:sec:appendix:strategies). In terms of Hypervolume, all five strategies outperform the “Retrain” approach, with the “Pre&Post” strategy achieving the best results. For Pearson r-a and Pearson r-d, the “Adap-norm” strategy demonstrates the best overall performance, indicating strong consistency with the “Retrain” approach, i.e., high controllability. Additionally, the Hypervolume remains within an acceptable range, suggesting that adding conditions aggregated by an attention mechanism to the layer norm is a promising approach for controllability.
6 Related Work

### v2

@@ 4.2. Training of Parameter Generation Model
(8) $\ell_{\text{diff}}=\mathbb{E}_{\bm{\theta}_{i,0}^{\text{a}},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\bm{\xi}}(\bm{\theta}_{i,t}^{\text{a}},\bm{w}_{i},t)\right\|^{2}\right],$
where $\epsilon$ denotes the noise to obtain $\bm{\theta}_{i,t}^{\text{a}}$ from $\bm{\theta}_{i,0}^{\text{a}}$, and the denoising model $\epsilon(\cdot)$ is the main part of the generative model $g_{\bm{\xi}}$. We assume that the parameters of $g_{\bm{\xi}}$ primarily originate from the denoising model. For simplicity, we denote the denoising model as $\epsilon_{\bm{\xi}}$. To conduct condition training in a classifier-free guidance manner (Ho and Salimans, 2022), we use the denoising model to serve as both the conditional and unconditional model by simply inputting a null token $\varnothing$ as the condition (i.e., preference weights $\bm{w}_{i}$) for the unconditional model, i.e. $\epsilon_{\bm{\xi}}(\bm{\theta}_{i,t}^{\text{a}},t)=\epsilon_{\bm{\xi}}(\bm{\theta}_{i,t}^{\text{a}},\bm{w}_{i}=\varnothing,t)$. The probability of setting $\bm{w}_{i}$ to $\varnothing$ is denoted as $p_{\mathrm{uncond}}$ and is configured as a hyperparameter.
4.3. Test-Time Parameter Generation
@@
(9) $\tilde{\epsilon}_{\bm{\xi}}(\bm{\theta}_{n,t}^{\text{a}},\bm{w}_{n},t)$ $=(1+\gamma)\epsilon_{\bm{\xi}}(\bm{\theta}_{n,t}^{\text{a}},\bm{w}_{n},t)-\gamma\epsilon_{\bm{\xi}}(\bm{\theta}_{n,t}^{\text{a}},t),$
@@
where $\bm{z}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ for $t>1$ and $\bm{z}_{t}=\bm{0}$ for $t=1$, $\beta_{t}=1-\alpha_{t}$, $\gamma\in[0,1]$ .
@@
Table 1. Statistical information of Datasets
Dataset #user #item #category #inter. density
MovieLens 1M 6,034 3,125 18 994,338 7.805%
Amazon Food 4,905 2,420 156 53,258 0.448%
Industrial Data 3,628 3,181 27 72,391 0.627%
5. Experiments
We conducted experiments to evaluate the performance of Paragon.
5.1. Experiment Settings
Table 2. Performance comparison between the proposed method and baseline models. The best results are highlighted in bold, while the second-best results are underlined. “/” represents the absence of a relevant value.
MovieLens Amazon Food Industrial Dataset
Backbone Algorithm Avg.HV pearson r-a pearson r-d Avg.HV pearson r-a pearson r-d Avg.HV pearson r-a pearson r-d
Retrain 0.2281 / / 0.2251 / / 0.2779 / /
CMR 0.1920 0.8901 0.9150 0.1955 -0.7039 0.9932 0.2476 0.8750 0.9237
Soup 0.1441 0.7861 0.9133 0.1561 0.5317 0.6693 0.1825 0.7306 0.8188
MMR 0.1808 0.9575 0.8803 0.1707 0.1320 -0.3087 0.2034 0.9077 0.9655
SASRec Paragon 0.2138 0.9905 0.9903 0.2420 0.8857 0.9816 0.2812 0.9976 0.9986
/ LLM-CMTR 0.0625 -0.0600 0.0994 0.1017 0.7296 0.8558 0.0372 0.7279 0.7132
Retrain 0.1823 / / 0.1556 / / 0.1735 / /
CMR 0.1760 0.9068 0.8813 0.3617 0.8287 0.9059 0.1230 0.6514 0.5985
Soup 0.1197 0.8061 0.6694 0.0604 0.3850 0.8005 0.1226 0.7099 0.8200
MMR 0.1609 0.8692 0.7257 0.1354 -0.3497 -0.3748 0.1287 0.7916 0.7553
GRU4Rec Paragon 0.2009 0.9929 0.9786 0.1623 0.8470 0.9685 0.1871 0.9760 0.9236
/ LLM-CMTR 0.0625 -0.0716 0.0119 0.0667 0.7484 0.7904 0.0372 0.8088 0.7722
Retrain 0.2301 / / 0.2232 / / 0.2777 / /
CMR 0.1769 0.9286 0.9903 0.2064 -0.6279 0.9828 0.3315 0.8855 0.9300
Soup 0.1483 0.8033 0.8858 0.1533 0.5342 0.6525 0.1811 0.7310 0.8248
MMR 0.1815 0.8946 0.8684 0.1672 0.3057 0.2037 0.2060 0.9004 0.9565
TiSASRec Paragon 0.2532 0.9923 0.9914 0.2394 0.8759 0.9851 0.2862 0.9968 0.9984
/ LLM-CMTR 0.0625 -0.0663 0.0999 0.0667 0.7213 0.8499 0.0372 0.7373 0.7451
5.1.1. Dataset
The datasets are processed as follows:
MovieLens-1M  [^1] https://grouplens.org/datasets/movielens/ is an website dataset of MovieLens in 2000. We sorted each user’s browsing history chronologically and filtered out users with fewer than 5 interactions. Each interaction is formatted to include user ID, item ID, timestamp, Categories of items (may be multiple categories).
Amazon Grocery and Gourmet Food  [^2] http://jmcauley.ucsd.edu/data/amazon/links.html is the food data from Amazon website, spanning from August 09, 2000 to July 23, 2014. Since the items belong to 156 categories, we used the GloVe (Pennington et al., 2014) to generate embeddings for each category. We then applied K-means clustering to group them into 30 broader categories. The interaction format is the same as MovieLens 1M.
The industrial dataset is the user click dataset from an electronics commercial store, spanning from July 24, 2024, to August 24, 2024. We randomly sample 10000 user, then filtering out users with fewer than 10 interactions to obtain 3628 users. Each interaction was formatted to match the structure used in MovieLens 1M. The specific statistical information of the three datasets is in Table 1.
5.1.2. Baselines
To validate its effectiveness, we compare our model against several baseline methods adapted for CMTR.
Retraining is performed using Linear Scalarization (Birge and Louveaux, 2011) based on task requirements, representing the optimal solution without considering the associated costs.
Soup (Wortsman et al., 2022) trains separate models for each aspect of the task requirement and merges them linearly during the testing stage.
MMR (Carbonell and Goldstein, 1998) is a heuristic post-processing approach with the item selected sequentially according to maximal marginal relevance.
CMR (Chen et al., 2023b) dynamically adjusts models based on preference weights using policy hypernetworks to generate model parameters.
LLM-CMTR (Dai et al., 2023) is a prompt-based method but specifically customized for CMTR. It inputs prompts containing specific preference weights to guide the LLM’s generation in the form of list-wise recommendations. In our experiments, we selected the llama3-7B-Instruct model.
5.1.3. Implement Details
We first filtered for 5-core user IDs and adopted a leave-one-out data splitting strategy to divide the dataset into training, validation, and test sets. Additionally, we limited the length of the target item’s interaction history to no more than 20. For the three recommendation model backbones, we follow the settings used in RecChorus, where the Adam optimizer is applied with a learning rate of 1e-3, embedding size of 64, and hidden size of 64. The number of negative samples in the training set is set to 9, and 99 for both the validation and test sets.
5.1.4. Metrics
@@
We evaluate the algorithms from two dimensions. Specifically, we use Hypervolume (HV) (Guerreiro et al., 2021) to measure the performance of the algorithm on each task, particularly in terms of the trade-offs between accuracy and diversity. The average HV (denoted as Avg.HV) across multiple tasks is used to assess the overall performance of the algorithm in balancing both objectives (accuracy and diversity). To eliminate the differences in scale between the two objectives, we normalize the performance on each objective. Additionally, we utilize the Pearson correlation coefficient to evaluate the alignment between the algorithm’s performance across different tasks and the optimal model, providing insight into the algorithm’s controllability. Pearson r-a, Pearson r-d measure the correlation between the algorithm and the optimal in terms of accuracy and diversity. Notably, the accuracy and diversity is measured by NDCG@10@10 and $\alpha$-NDCG@10). The fairness in analysis experiment is measured by Absolute Difference (AD) (Wang et al., 2023b).
@@
(b) Diversity on MovieLens 1M
(c) Accuracy on Amazon Food
(d) Diversity on Amazon Food
(e) Accuracy on Industrial Data
@@
Figure 4. The accuracy and diversity curve of Paragon and other baselines in NDCG@10 and $\alpha$-NDCG@10 across accuracy weights ranging from 0 to 1, with intervals of 0.1. The backbone is TiSASRec.
5.2. Experimental Results
We conducted experiments to address the following two questions: i) How transferable is Paragon, specifically in terms of its ability to adapt to different backbone algorithms? ii) How does Paragon perform compared to other baselines? The results are presented in Table 2.
To answer the first question, we used commonly adopted sequential recommendation models as backbones (e.g., SASRec (Kang and McAuley, 2018), GRU4Rec (Hidasi, 2015), and TiSASRec (Li et al., 2020b)) and conducted extensive experiments across three datasets. Specifically, we evaluated Paragon’s performance under various task descriptions by measuring NDCG@10 and $\alpha$-NDCG@10. The accuracy weight $w_{\text{acc.}}$ varies from 0 to 1 in intervals of 0.1, with the corresponding diversity weight set as $w_{\text{div.}}=1-w_{\text{acc.}}$. We then post-processed NDCG@10 and $\alpha$-NDCG@10 across different tasks to compute Avg.HV, Pearson r-a, and Pearson r-d. These metrics respectively evaluate the quality of multi-objective optimization on individual tasks and the controllability across multiple tasks.
Overall, across the three backbones and three datasets, Paragon consistently ranked among the top two performers across all three evaluation metrics. Notably, in most cases, the top two of Avg.HV are Retrain and Paragon, indicating that Paragon’s performance in multi-objective trade-offs is on par with, or even superior to, Retrain. Specific exceptions occurred, such as on the Amazon Food dataset with GRU4Rec as the backbone and the Industrial Data with TiSASRec as the backbone, where CMR achieved the best Avg.HV. This is because CMR is not influenced by task descriptions and thus maintains consistently high NDCG@10 scores (as shown in Figure 4 and further explained in response to question ii). For Pearson r-a and Pearson r-d, Paragon demonstrated strong correlations with the Retrain method, indicating that Paragon closely aligns with Retrain (which we assume to be optima) in terms of accuracy (NDCG@10) and diversity ($\alpha$-NDCG@10) across different tasks. On the Amazon Food dataset with SASRec as the backbone, CMR achieved the highest Pearson r-d. However, its Pearson r-a was negative, indicating a lack of control and a collapse in accuracy.
@@ 5.2. Experimental Results
Figure 5. The variation of performance before and after disturbance in MovieLens 1M based on SASRec. The blue bars represent the variation in NDCG@10, while the red represent the variation in $\alpha$-NDCG@10
To address the second question, we presented the specific performance of Paragon under each task description using TiSASRec as the backbone across three datasets, as shown in Figure 4. It is observed that in all three datasets, Paragon’s NDCG@10 progressively increases with higher accuracy weights, while $\alpha$-NDCG@10 decreases correspondingly due to the simultaneous reduction in diversity weight. These trends demonstrate the effectiveness of our algorithm in controllability. Notably, assuming that Retrain is optimal, Paragon exhibits strong consistency with the Retrain method. In contrast, MMR, as a post-processing algorithm, shows variability because varying degrees of diversity manipulation can disrupt the original recommendation list, uncontrollably affecting its accuracy. The Soup method merges the parameters of accuracy and diversity models based on their weights, aligning closely with the Retrain model when accuracy weights are extreme but showing significant deviations in other tasks. This indicates that tasks do not follow a simple linear relationship with different preference weights, and Soup makes overly strong assumptions about this relationship. CMR demonstrates inconsistent performance across different datasets. On the MovieLens 1M dataset, CMR aligns well with the original descriptions by exhibiting high diversity. However, on the other two datasets, it shows a stable yet uncontrollable state; for instance, on the Amazon Food dataset, CMR maintains high accuracy even with low accuracy weights, and on the Industrial Data, it retains high diversity despite low diversity weights.
(a) The result of fairness.
(b) The result of acc. and div.
Figure 6. The result of control under accuracy, diversity and fairness. The accuracy weight ranges from 0 to 1, with intervals of 0.1. The diversity weight is set to 11- accuracy weight. The fairness weight is $\{1,0\}$ corresponding to ‘fair’ and ‘unfair’. The experiment is conducted on the MovieLens 1M, uti-lizing SASRec as the backbone.
5.3. Analyses
We conducted our analysis experiments based on three key research questions:
Table 3. Response time comparison between proposed Paragon and the “Retrain” approach across three datasets using three different backbones. Note that the unit is seconds (sec.).
Approach Backbone MovieLens 1M (sec.) Amazon Food (sec.) Industrial Data (sec.)
Retrain SASRec 293.10 ± 11.61 91.01 ± 2.34 46.82 ± 3.25
GRU4Rec 281.60 ± 17.36 92.39 ± 4.28 49.54 ± 2.38
TiSASRec 303.80 ± 9.09 105.40 ± 7.66 52.47 ± 4.64
Paragon SASRec 2.68 ± 0.36-99.1% 2.64 ± 0.36-97.1% 2.55 ± 0.25-94.6%
GRU4Rec 2.56 ± 0.27-99.1% 2.54 ± 0.24-97.3% 2.51 ± 0.23 -94.9%
TiSASRec 2.55 ± 0.23-99.2% 2.52 ± 0.24-97.6% 2.58 ± 0.26 -95.1%
5.3.1. RQ1: What are the advantages of Diffusion over Hypernetwork in parameter generation?
We conducted experiments to validate the robustness of the parameters generated by Paragon. Specifically, we designed three sets of experiments to constructed adapter parameters: “Retrain”, “Diffusion” and “Hypernetwork”, where “Hypernetwork” utilizes the MLP to learn the relationship between the preference weight and the optimized adapter parameters. First, we add Gaussian noise of the same magnitude to all three sets of adapter parameters and measured the resulting fluctuations in NDCG@10 and $\alpha$-NDCG@10. We observed that the parameters generated by Paragon exhibited the lowest performance fluctuations than others, both in terms of accuracy (NDCG@10) and diversity ($\alpha$-NDCG@10) as depicted in Figure 5. We could draw the following conclusions: 1) the Diffusion approach demonstrates superior robustness in parameter generation across both accuracy and diversity metrics compared to alternative methodologies. Second, Hypernetwork-based methods exhibit inferior robustness in parameter generation, a phenomenon likely attributable to the inherent limitations of discriminative models in achieving optimal generative capacity.
(a) Pre.
(b) Pre.&Post.
(c) Pre-Adaptive
(d) Post- Adaptive
(e) Adaptive- Norm
Figure 7. Illustration of the structural diagram for five conditioning strategies.
5.3.2. RQ2: Is Paragon efficient enough to handle real-time changes in preference weights compared to Retrain?